# Lifting Architectural Constraints of Injective Flows

### v3

@@ 1 Introduction
Generative modeling is one of the most important tasks in machine learning, having numerous applications across vision (Rombach et al., 2022), language modeling (Brown et al., 2020), science (Ardizzone et al., 2018; Radev et al., 2020) and beyond. One of the best-motivated approaches to generative modeling is maximum likelihood training, due to its favorable statistical properties (Hastie et al., 2009). In the continuous setting, exact maximum likelihood training is most commonly achieved by normalizing flows (Rezende & Mohamed, 2015; Dinh et al., 2014; Kobyzev et al., 2020) which parameterize an exactly invertible function with a tractable change of variables (log-determinant term). This generally introduces a trade-off between model expressivity and computational cost, where the cheapest networks to train and sample from, such as coupling block architectures, require very specifically constructed functions which may limit expressivity (Draxler et al., 2022). In addition, normalizing flows preserve the dimensionality of the inputs, requiring a latent space of the same dimension as the data space.
Due to the manifold hypothesis (Bengio et al., 2013), which suggests that realistic data lies on a low-dimensional manifold embedded into a high-dimensional data space, it is more efficient to only model distributions on a low-dimensional manifold and regard deviations from the manifold as uninformative noise. Prior works such as Caterini et al. (2021); Brehmer & Cranmer (2020) have restricted normalizing flows to low-dimensional manifolds via specially-constructed bottleneck architectures (known as “invertible autoencoders” (Teng & Choromanska, 2019) or “injective flows” (Kothari et al., 2021)) where encoder and decoder share parameters. These injective flows are optimized by some version of maximum likelihood training. This is not an ideal design decision, as the restrictive architectures used in such models were originally introduced for tractable change of variables calculations in normalizing flows, but such calculations are not possible in the presence of a bottleneck (Brehmer & Cranmer, 2020). As a result, we propose to drop the restrictive constructions (such as coupling blocks), instead use an unconstrained encoder and decoder, and introduce a new technique to get around calculating the change of variables. This greatly simplifies the design of the model and makes it more expressive.
@@
We build on the unbiased maximum likelihood estimator used by rectangular flows (Caterini et al., 2021) to approximate the gradient of the change of variables term. We simplify the estimator considerably by replacing iterative conjugate gradient, which can require iterating up to the number of latent dimensions for convergence, with an efficient single-step estimator. This is fast: a batch can be processed in about twice the time (or less) as an autoencoder trained on reconstruction loss only. In addition, we make a novel observation about injective flows: naively training with maximum likelihood is ill-defined due to the possibility of diverging curvature in the decoding function. To fix this problem, we propose a modification to our maximum likelihood estimator which counteracts the possibility of diverging curvature. We call our model the free-form injective flow (FIF).
@@
  • We introduce an efficient maximum likelihood estimator for free-form injective flows, lifting architectural constraints on the models (section 4.1).
@@
  • We demonstrate competitive performance with a series of experiments on toy, tabular and image data (section 5).
@@ 2 Related work
One set of methods attempts to use full-dimensional normalizing flows, with some additional regularization or architectural constraints such that a subspace of the latent space corresponds to the manifold. One strategy adds noise to the data to make it a full-dimensional distribution then denoises to the manifold (Horvat & Pfister, 2021; Loaiza-Ganem et al., 2023). Another restricts the non-manifold latent dimensions to have small variance (Beitler et al., 2021; Silvestri et al., 2023; Zhang et al., 2023).
@@
Several works attempt to reduce the constraints imposed by typical normalizing flow architectures, allowing the use of free-form networks. However, all of these methods only apply to full-dimensional architectures. FFJORD (Grathwohl et al., 2018) is a type of continuous normalizing flow (Chen et al., 2018b) which estimates the change of variables stochastically. Residual flows (Behrmann et al., 2019; Chen et al., 2019) make residual networks invertible, but require expensive iterative estimators to train via maximum likelihood. Self-normalizing flows (Keller et al., 2021) and relative gradient optimization (Gresele et al., 2020) estimate maximum likelihood gradients for the matrices used in neural networks, but restrict the architecture to use exclusively square weight matrices without skip connections.
@@
Many methods optimize some bound on the full-dimensional maximum likelihood, notably the variational autoencoder (Kingma & Welling, 2013) and its variants. Cunningham et al. (2020) also optimizes a variational lower bound to the likelihood. Other methods fit into the injective flow framework by jointly optimizing a reconstruction loss and some approximation to maximum likelihood on the manifold: Kumar et al. (2020); Zhang et al. (2020) approximate the log-determinant of the Jacobian by its Frobenius norm. The entropic AE (Ghose et al., 2020) maximizes the entropy of the latent distribution by a nearest-neighbor estimator while constraining its variance, resulting in a Gaussian latent space. In addition, there are other ways to regularize the latent space of an autoencoder which are not based on maximum likelihood, e.g. adversarial methods (Makhzani et al., 2015).
@@ 3 Background
Let $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{d}$ be an encoder which compresses data to a latent space and a decoder $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ which decompresses the latent representation. A full-dimensional model has $d=D$ while a bottleneck model has $d<D$. If $f\circ g$ is the identity, then we call $f$ and $g$ consistent. For example, the forward and reverse function of a typical normalizing flow are consistent as $f^{-1}=g$.
@@
Normalizing flow models, trained by maximum likelihood via the change of variables theorem, are only well-defined when mapping between spaces of equal dimension. A result from differential geometry (Krantz & Parks, 2008) allows us to generalize the change of variables theorem to non-equal-dimension transformations through the formula:
@@
where $f$ and $g$ are consistent and primes denote derivatives: $g^{\prime}(f(x))$ is the Jacobian matrix of $g$ evaluated at $f(x)$. Note that, since $p_{X}$ is derived as the pushforward of the latent distribution $p_{Z}$ by $g$, the formula is valid only for $x$ which lie on the decoder manifold (see appendix A for more details). Unlike in the full-dimensional case, there is no architecture known which can represent arbitrary manifolds and allows efficient exact computation of eq. 2 (see section 2).
@@
The log-determinant term is the difficult one to optimize. Fortunately, its gradient with respect to the parameters $\theta$ of the decoder can be estimated tractably (Caterini et al., 2021). Note that $g=g_{\theta}$ but the $\theta$ subscript is dropped to avoid clutter. The relevant quantity is (with $J=g^{\prime}(z)$):
@@
The trace can be estimated (Hutchinson, 1989; Girard, 1989) as:
$\frac{\partial}{\partial\theta_{j}}\frac{1}{2}\log\det(J^{\top}J)\approx\frac{1}{2K}\sum_{k=1}^{K}\epsilon_{k}^{\top}(J^{\top}J)^{-1}\frac{\partial}{\partial\theta_{j}}(J^{\top}J)\epsilon_{k},$ (5)
where the $\epsilon_{k}$ are sampled from a distribution where $E[\epsilon\epsilon^{\top}]=\mathbb{I}$, typically either Rademacher or standard normal. Matrix multiplication can be avoided by using vector-matrix and matrix-vector product subroutines. These are readily obtained by automatic differentiation in the case of vector products with $J^{\top}J$ and by autodiff combined with the conjugate gradient iterative method for $(J^{\top}J)^{-1}$ products. Write $\epsilon_{k}^{\top}(J^{\top}J)^{-1}=\texttt{CG}(J^{\top}J;\epsilon_{k})^{\top}$ where $\texttt{CG}(A;b)$ denotes the solution to $Ax=b$ via conjugate gradient. The parameter derivative can be made to act only on the rightmost Jacobian terms by applying the stop-gradient operation, to the output of the conjugate gradient method. The final surrogate loss function for the log-determinant term is therefore:
@@
which replaces the log-determinant term in the loss. The function stop_gradient returns its input, but makes it constant with respect to parameters. Note that the resulting loss does not have the same value as the original loss in eq. 3, but has the same gradient. In the following, we present a significantly improved version of this estimator based on the insight that the encoder is an (approximate) inverse of the decoder.
@@ 4 Free-form injective flow (FIF)
with $z=f(x)$. Note the negative sign before the surrogate term, which comes from sending the log-determinant gradient to the encoder rather than the decoder. We will derive and motivate this formulation of the loss in sections 4.1, 4.2 and 4.2.
@@ 4.1 Simplifying the surrogate
We considerably simplify the optimization of rectangular flows by a new surrogate for the log-determinant term, which uses the Jacobian of the encoder as an approximation for the inverse Jacobian of the decoder. This allows the surrogate to be computed in a single pass, avoiding costly conjugate gradient iterations.
@@
where $J^{\dagger}=(J^{\top}J)^{-1}J^{\top}$ is the Moore-Penrose inverse of $J$. The full derivation is in appendix B. To see the advantage of this formulation consider that for $f$ and $g$ optimal with respect to the reconstruction loss, $f^{\prime}(\hat{x})=g^{\prime}(f(x))^{\dagger}$ (see section B.1). Using the stop_gradient operation, this leads to the following surrogate loss term:
@@
Each term of the sum can be computed from just two vector-Jacobian products obtained from backward-mode automatic differentiation. This is a significant improvement on the iterative conjugate gradient method needed in the original formulation of rectangular flows which requires up to $2(d+1)$ vector-Jacobian or Jacobian-vector products to ensure convergence (Caterini et al., 2021). We measure 1.5$\sim$ 1.5$\times$ 1.5 × to 2×2$\times$2 × the wall clock time of our loss compared to reconstruction loss only, independent of the latent dimension. Note that the surrogate is only accurate if $f$ and $g$ are (at least approximately) optimal with respect to the reconstruction loss. We assume that this is sufficiently fulfilled in our model since we directly optimize the reconstruction loss and we observe stable training in practice, thus validating the assumption.
@@ 4.2 Problems with maximum likelihood in the presence of a bottleneck
Rectangular flows are trained with a combination of a reconstruction and likelihood term. We might ask what happens if we only train with the likelihood term, making an analogy to normalizing flows. In this case our loss would be:
@@
Unfortunately, optimizing this loss can lead us to learn a degenerate decoder manifold, an issue raised in Brehmer & Cranmer (2020). Here we expand on their argument to show why the model will tend to learn a manifold which aligns with the lowest entropy directions of the data. We show how to fix this problem in the linear case with the addition of a reconstruction term but show that adding a reconstruction term in the nonlinear case is not sufficient to avoid pathological solutions.
First consider that the per-sample loss is invariant to projections: $\mathcal{L}_{\textrm{NLL}}(\hat{x})=\mathcal{L}_{\textrm{NLL}}(x)$, since $\mathcal{L}_{\textrm{NLL}}(x)$ is a function only of $f(x)$ and $f(\hat{x})=f(x)$. This means that we can write our loss as:
@@
$KL(\hat{p}_{\textrm{data}}(\hat{x})||p_{\theta}(\hat{x}))=-H(\hat{p}_{\textrm{data}}(\hat{x}))-E_{p_{\textrm{data}}(\hat{x})}[\log p_{\theta}(\hat{x})]\geq 0.$ (13)
@@
Unlike in standard normalizing flow optimization, where the right hand side would be fixed, here the entropy on the right depends on the learned model, and the loss can continue to decrease without bound by reducing the entropy of the projected data. See fig. 2 (left) for an illustration of a pathological case. This is a toy model with a 1-dimensional latent space where the decoder maps to a semicircle with learnable radius which is offset horizontally. The encoder projects orthogonally onto the manifold. Naively optimizing the likelihood leads to increasing curvature since this decreases the entropy of the projected data indefinitely.
@@
Since it is clear that optimization of a likelihood loss without reconstruction will fail, we ask what happens as we reintroduce a reconstruction loss at various strengths. Suppose the loss is of the form:
$\mathcal{L}=\mathcal{L}_{\textrm{NLL}}+\beta E_{p_{\textrm{data}}(x)}\left[\lVert\hat{x}-x\rVert^{2}\right].$ (15)
In appendix C the closed-form solution to this model is worked out in detail when $f$ and $g$ are linear functions. The model learns low-entropy directions of the data when $\beta$ is too low but transitions to the PCA solution for large enough $\beta$. This transition occurs at $\beta=1/2\sigma^{2}$ where $\sigma$ is the smallest eigenvalue of the data covariance matrix.
Unfortunately, when $f$ and $g$ are non-linear, the addition of a reconstruction term is not enough to fix the pathological behavior of the likelihood loss defined on the manifold. Consider again fig. 2, where the left-hand figure is optimized with both likelihood and reconstruction loss. Without additional constraints on the decoding function, the curvature can increase without bound, leading to a loss tending towards negative infinity.
@@
The term which leads to pathological behavior in the likelihood loss is the log-determinant. We make the fairly simple modification of evaluating $f^{\prime}$ in our estimator at $x$ rather than $\hat{x}$. Namely, we modify eq. 10 to estimate the gradient of the log-determinant term by:
@@
When using the change of variables with $f^{\prime}$ evaluated at $\hat{x}$, all that matters is the change of volume from the projected data to the latent space, so we can decrease the loss by choosing a manifold which concentrates the projected data more tightly (the more possibility it has to expand the data, the lower the loss will be). We can counteract this effect by introducing a factor inversely proportional to the concentration, leading to the stated correction. See section B.2 for a detailed explanation.
In this way, we discourage pathological solutions involving high curvature. In fig. 2 (right) we can see the effect of the modified estimator: the manifold now moves towards the data since the optimization is not dominated by diverging curvature.
@@
Figure 3 shows that when using this loss, if $\beta$ is large enough, the dominant manifold direction is identified. In section B.1, we show a similar experiment on MNIST.

### v4

@@ 1 Introduction
Generative modeling is one of the most important tasks in machine learning, having numerous applications across vision (Rombach et al., 2022), language modeling (Brown et al., 2020), science (Ardizzone et al., 2018; Radev et al., 2022) and beyond. One of the best-motivated approaches to generative modeling is maximum likelihood training, due to its favorable statistical properties (Hastie et al., 2009). In the continuous setting, exact maximum likelihood training is most commonly achieved by normalizing flows (Rezende & Mohamed, 2015; Dinh et al., 2015; Kobyzev et al., 2021) which parameterize an exactly invertible function with a tractable change of variables (log-determinant term). This generally introduces a trade-off between model expressivity and computational cost, where the cheapest networks to train and sample from, such as coupling block architectures, require very specifically constructed functions which may limit expressivity (Draxler et al., 2022). In addition, normalizing flows preserve the dimensionality of the inputs, requiring a latent space of the same dimension as the data space.
Due to the manifold hypothesis (Bengio et al., 2013), which suggests that realistic data lies on a low-dimensional manifold embedded into a high-dimensional data space, it is more efficient to only model distributions on a low-dimensional manifold and regard deviations from the manifold as uninformative noise. Prior works such as Caterini et al. (2021); Brehmer & Cranmer (2020) have restricted normalizing flows to low-dimensional manifolds via specially-constructed bottleneck architectures (known as “invertible autoencoders” (Teng & Choromanska, 2019) or “injective flows” (Kothari et al., 2021)) where encoder and decoder share parameters. These injective flows are optimized by some version of maximum likelihood training. This is not an ideal design decision, as the restrictive architectures used in such models were originally designed for tractable change of variables calculations in normalizing flows, but such calculations are not possible in the presence of a bottleneck (Brehmer & Cranmer, 2020). As a result, we propose to drop the restrictive constructions (such as coupling blocks), instead use an unconstrained encoder and decoder, and introduce a new technique to get around calculating the change of variables. This greatly simplifies the design of the model and makes it more expressive.
@@
We build on the unbiased maximum likelihood estimator used by rectangular flows (Caterini et al., 2021) to approximate the gradient of the change of variables term. We simplify the estimator considerably by replacing iterative conjugate gradient with an efficient single-step estimator. This is fast: a batch can be processed in about twice the time (or less) as an autoencoder trained on reconstruction loss only. In addition, we make a novel observation about injective flows: naively training with maximum likelihood is ill-defined due to the possibility of diverging curvature in the decoding function. To fix this problem, we propose a modification to our maximum likelihood estimator which counteracts the possibility of diverging curvature. We call our model the free-form injective flow (FIF).
@@
  • We introduce an efficient maximum likelihood estimator for free-form injective flows and use it to train an unconstrained injective flow for the first time (section 4.1).
@@
  • We outperform previous injective flows and demonstrate competitive performance to generative autoencoders on toy, tabular and image data (section 5). We provide code to implement our model and reproduce our results at https://github.com/vislearn/FFF.
@@ 2 Related work
One set of methods attempts to use full-dimensional normalizing flows, with some additional regularization or architectural constraints such that a subspace of the latent space corresponds to the manifold. One strategy adds noise to the data to make it a full-dimensional distribution then denoises to the manifold (Horvat & Pfister, 2021; Loaiza-Ganem et al., 2022). Another restricts the non-manifold latent dimensions to have small variance (Beitler et al., 2021; Silvestri et al., 2023; Zhang et al., 2023).
@@
Several works attempt to reduce the constraints imposed by typical normalizing flow architectures, allowing the use of free-form networks. However, all of these methods only apply to full-dimensional architectures. FFJORD (Grathwohl et al., 2019) is a type of continuous normalizing flow (Chen et al., 2018b) which estimates the change of variables stochastically. Residual flows (Behrmann et al., 2019; Chen et al., 2019) make residual networks invertible, but require expensive iterative estimators to train via maximum likelihood. Self-normalizing flows (Keller et al., 2021) and relative gradient optimization (Gresele et al., 2020) estimate maximum likelihood gradients for the matrices used in neural networks, but restrict the architecture to use exclusively square weight matrices without skip connections. In a parallel work, we present the extension of the present paper to full-dimensional normalizing flows based on free-form neural networks, called free-form flows (Draxler et al., 2024).
@@
Many methods optimize some bound on the full-dimensional maximum likelihood, notably the variational autoencoder (Kingma & Welling, 2014) and its variants. Cunningham et al. (2020) also optimizes a variational lower bound to the likelihood. Other methods fit into the injective flow framework by jointly optimizing a reconstruction loss and some approximation to maximum likelihood on the manifold: Kumar et al. (2020); Zhang et al. (2020) approximate the log-determinant of the Jacobian by its Frobenius norm. The entropic AE (Ghose et al., 2020) maximizes the entropy of the latent distribution by a nearest-neighbor estimator while constraining its variance, resulting in a Gaussian latent space. In addition, there are other ways to regularize the latent space of an autoencoder which are not based on maximum likelihood, e.g. adversarial methods (Makhzani et al., 2015).
@@ 3 Background
Let $f:\mathbb{R}^{D}\rightarrow\mathbb{R}^{d}$ be an encoder which compresses data to a latent space and a decoder $g:\mathbb{R}^{d}\rightarrow\mathbb{R}^{D}$ which decompresses the latent representation. A full-dimensional model has $d=D$ while a bottleneck model has $d<D$. If $f\circ g:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the identity, then we call $f$ and $g$ consistent. For example, the forward and inverse function of a normalizing flow are consistent as $f^{-1}=g$.
@@
The maximum likelihood objective resulting from the change of variables theorem, used to train normalizing flow models, is only well-defined when mapping between spaces of equal dimension. A result from differential geometry (Krantz & Parks, 2008) allows us to generalize the change of variables theorem to non-equal-dimension transformations through the formula:
@@
where $f$ and $g$ are consistent and primes denote derivatives: $g^{\prime}(f(x))$ is the Jacobian matrix of $g$ evaluated at $f(x)$. Note that, since $p_{X}$ is derived as the pushforward of the latent distribution $p_{Z}$ by $g$, the formula is valid only for $x$ which lie on the decoder manifold (see appendix A for more details). Unlike in the full-dimensional case, there is no architecture known which can represent arbitrary manifolds and at the same time allows efficient exact computation of eq. 2 (see section 2).
@@
The log-determinant term is the difficult one to optimize. Fortunately, its gradient with respect to the parameters $\theta$ of the decoder can be estimated tractably by the following construction (Caterini et al., 2021). Note that $g=g_{\theta}$ but the $\theta$ subscript is dropped to avoid clutter. The relevant quantity is (with $J=g^{\prime}(z)$):
@@
The trace can be estimated (Hutchinson, 1989; Girard, 1989) by:
$\frac{\partial}{\partial\theta_{j}}\frac{1}{2}\log\det(J^{\top}J)\approx\frac{1}{2K}\sum_{k=1}^{K}\underbrace{\epsilon_{k}^{\top}(J^{\top}J)^{-1}}_{\text{($**$)}}\frac{\partial}{\partial\theta_{j}}\underbrace{(J^{\top}J)\epsilon_{k}}_{\text{($*$)}},$ (5)
where the $\epsilon_{k}$ are $K$ samples from a distribution where $E[\epsilon\epsilon^{\top}]=\mathbb{I}$, typically either Rademacher or standard normal. Now, we take steps to compute the above without constructing the entire network Jacobian $J$. First, note that (*) can be written as a Jacobian-vector product $v_{1}=Jv$ and a vector-Jacobian product $J^{T}v_{1}=(v_{1}^{\top}J)^{\top}$, each readily available via automatic differentiation. For (**∗ ∗), Caterini et al. (2021) propose to employ the iterative conjugate gradient method: Write $\epsilon_{k}^{\top}(J^{\top}J)^{-1}=\texttt{CG}(J^{\top}J;\epsilon_{k})^{\top}$ where $\texttt{CG}(A;b)$ denotes the conjugate gradient solution to $Ax=b$. Conjugate gradient is a clever choice here, since it again only requires computing terms of the form $J^{\top}Jv$ using autodiff. The parameter derivative can be made to act only on the rightmost Jacobian terms by applying the stop_gradient operation to the output of the conjugate gradient method, which returns its input, but has zero gradient. The final surrogate for the log-determinant term is therefore:
@@
which replaces the log-determinant term in the loss. This computation yields the same gradient as the original loss in eq. 3, even though it has a different value. In the following, we present a significantly improved version of this estimator based on the insight that the encoder is an (approximate) inverse of the decoder.
@@ 4 Free-form injective flow (FIF)
with $z=f(x)$. Note the negative sign before the surrogate term, which comes from sending the log-determinant gradient to the encoder rather than the decoder. We will derive and motivate this formulation of the loss in sections 4.1 and 4.2.
@@ 4.1 Simplifying the surrogate
We considerably simplify the optimization of rectangular flows by a new surrogate for the log-determinant term in eq. 2, which uses the Jacobian of the encoder as an approximation for the inverse Jacobian of the decoder. This allows the surrogate to be computed in a single pass, avoiding costly conjugate gradient iterations.
@@
where $J^{\dagger}=(J^{\top}J)^{-1}J^{\top}$ is the Moore-Penrose inverse of $J$. The full derivation is in appendix B. To see the advantage of this formulation consider that for encoder $f$ and decoder $g$ optimal with respect to the reconstruction loss, $f^{\prime}(\hat{x})=g^{\prime}(f(x))^{\dagger}$ (see section B.1). Using the stop_gradient operation, this leads to the following surrogate loss term:
@@
Each term of the sum can be computed from just two vector-Jacobian/Jacobian-vector products obtained from automatic differentiation. This is a significant improvement on the iterative conjugate gradient method needed in the original formulation of rectangular flows which requires up to $2(d+1)$ such products to ensure convergence (Caterini et al., 2021). Compared to reconstruction loss only, we measure 1.5$\sim$ 1.5$\times$ 1.5 × to 2×2$\times$2 × the wall clock time, independent of the latent dimension. Note that the surrogate is only accurate if $f$ and $g$ are (at least approximately) optimal with respect to the reconstruction loss. We observe stable training in practice, validating this assumption.
@@ 4.2 Problems with maximum likelihood in the presence of a bottleneck
Rectangular flows are trained with a combination of a reconstruction and a likelihood term. We might ask what happens if we only train with the likelihood term, making an analogy to normalizing flows. In this case our loss would be:
@@
Unfortunately, optimizing this loss can lead us to learn a degenerate decoder manifold, an issue raised in Brehmer & Cranmer (2020). Here we expand on their argument. First consider that if $f$ and $g$ are consistent, then $f(\hat{x})=f(g(f(x)))=f(x)$ and the per-sample loss is invariant to projections: $\mathcal{L}_{\textrm{NLL}}(\hat{x})=\mathcal{L}_{\textrm{NLL}}(x)$, since $\mathcal{L}_{\textrm{NLL}}(x)$ is a function only of $f(x)$ This means that we can write our loss as:
@@
$\operatorname{KL}(\hat{p}_{\textrm{data}}(\hat{x})\|p_{\theta}(\hat{x}))=-H(\hat{p}_{\textrm{data}}(\hat{x}))-E_{\hat{p}_{\textrm{data}}(\hat{x})}[\log p_{\theta}(\hat{x})]\geq 0.$ (13)
@@
Unlike in standard normalizing flow optimization, where the right hand side would be fixed, here the entropy depends on the projection learned by the model. Thus, the model could modify the projection such that entropy is as low as possible. We break this pathology down into two cases:
  1. 1.
    A model manifold which does not align with the data manifold but instead intersects it. For example, Brehmer & Cranmer (2020) discuss a case where a linear model learns to project a data distribution to a single point on the manifold, thus reducing its entropy to $\infty$- ∞, the lowest possible value. To the best of our knowledge, this can be fixed by adding noise and a reconstruction loss with sufficiently high weight. In appendix C we prove as much for linear models and characterize the solutions, which are the same as PCA if $\beta\geq 1/2\sigma^{2}$ where $\sigma^{2}$ is the smallest eigenvalue of the data covariance matrix.
  2. 2.
    A model manifold which concentrates the data by use of high curvature, see fig. 2 (left). This newly identified pathological case only occurs in nonlinear models, hence Brehmer & Cranmer (2020) did not notice this effect in their linear example. Importantly, this is not fixed by adding a reconstruction loss.
Most existing injective flows avoid this by a two-stage training, which first learns a projection and then the distribution of the projected data in the latent space. To enable jointly learning a manifold and a maximum-likelihood density on it, we need to find a fix for the pathology.
@@
The term which leads to pathological behavior in the likelihood loss is the log-determinant. When using the change of variables with $f^{\prime}$ evaluated at $\hat{x}$, all that matters is the change of volume from the projected data to the latent space, so the model can decrease the loss by choosing a manifold which concentrates the projected data more tightly (the more possibility it has to expand the data, the lower the loss will be). We can counteract this effect by introducing a factor inversely proportional to the concentration. This can be achieved by the fairly simple modification of evaluating $f^{\prime}$ in our estimator at $x$ rather than $\hat{x}$. Namely, we modify eq. 10 to estimate the gradient of the log-determinant term by:
@@
See section B.2 for a detailed explanation.
In this way, we discourage pathological solutions involving high curvature. In fig. 2 (right) we can see the effect of the modified estimator: the manifold now moves towards the data since the optimization is not dominated by diverging curvature. We note that the modified estimator is also computationally cheaper, since the vector-Jacobian product $\epsilon_{k}^{\top}f^{\prime}(x)$ can reuse the computational graph generated when computing $z$.
@@
Figure 3 shows that when using this loss, if $\beta$ is large enough, the dominant manifold direction is identified. In section E.1.2, we show a similar experiment on MNIST.